Anderson​ Motors, Inc. has just set the company dividend policy at ​$0.90 per year. The company plans to be in business forever. What is the price of this stock if a.  an investor wants a return of 3​%? b.  an investor wants a return of 8​%? c.  an investor wants a return of 9​%? d.  an investor wants a return of 12​%? e.  an investor wants a return of 17​%?
Answer #1

Use the constant dividend infinite dividend stream model:

Price = Dividend / r

a. Price = $0.90 / 0.03 = $30

b. Price = $0.90 / 0.08 = $11.25

c. Price = $0.90 / 0.09 = $10

d.  Price = $0.90 / 0.12 = $7.50

e.  Price = $0.90 / 0.17 = $5.29